OpenStudy (anonymous):

Consider the functions f(x) = 2(x - 7) and g(x) = 2x. What is the value of (f + g)(4)? Answer -2 14 -6 2

OpenStudy (anonymous):

2

OpenStudy (jamesj):

There are two ways to tackle this problem. The first one is to notice that (f + g)(4) = f(4) + g(4) So if you find f(4) and g(4), you just need to add them together.

OpenStudy (jamesj):

The second way is to find the function (f+g)(x). Once you've found it, evaluate it for x = 4.

OpenStudy (jamesj):

So looking at the first way, if f(x) = 2(x - 7) and g(x) = 2x then f(4) = 2(4-7) = 2(-3) = -6 and g(4) = 8 Hence f(4) + g(4) = -6 + 8 = 2. ======== Now the second way, (f+g)(x) = 2(x-7) + 2x = 2x - 14 + 2x = 4x - 14 Hence (f+g)(4) = 4(4) - 14 = 16 - 14 = 2 the same answer, which is just as well!

OpenStudy (jamesj):

Make sense?